# Optimization of a cosmic muon tomography scanner for cargo border control inspection

### v1

@@ IV GEANT4 Study of Secondary Hits
As muons pass through the hodoscope, cargo, container and surrounding air they can produce secondary particles, primarily delta rays. An assessment is made in this section of the impact of any hodoscope hits induced by such secondary activity. Note that activity due to other particles generated in cosmic air showers is not considered in this study but can also be considered as a source of ‘secondary’ information in MST.
@@
The goal of these studies was to evaluate what impact deposits from secondaries could have in degrading the performance of MST with the SilentBorder scanners. It is clear from FIG. 7 that above a certain energy threshold there is no method to distinguish between secondary and muon hits based on the scale of the energy deposition.
For the first study a threshold of 0.4 MeV is applied, retaining only those depositions that lie within the MIP peak region. A cargo block of size 1 m1 m0.5 m is placed at the center of the container with a variety of material types considered. The MuSiBo generator is used to provide muons corresponding to a 15 minute exposure time. The events recorded are then filtered to only retain events in which at least two hits were recorded before and after the container. This normally corresponds to three hits in the top set of hodoscopes plus a further three in either the side or bottom hodoscopes, occasionally the first set of hits will be found in one of the side hodoscopes. The goal of this filtering is to select events in which particle trajectories before and after the cargo could be selected, the scattering angle reconstructed and MST performed. FIG. 8(a) shows the total number of hits recorded across all hodoscopes per event, regardless of whether they were produced via secondaries or muons. In general the peak structure around 6 hits is produced via the 3 hits left by the muon, before and after the cargo whilst the small tail is produced by the presence of any secondary hits. Whilst it appears that there is some variation between the different cargo types this variation stems from the filtering and combining of hits, not from a significant difference in the number of secondary hits produced per muon. Higher density and higher-Z materials reduce the number of events that pass the filtering process. FIG. 8(b) shows the same distribution but normalized by the number of events that passed the filtering, very limited variation per cargo is now observed.
@@
In conclusion, it is clear that with the current SilentBorder scanner setup the impact of secondary hits is very limited. The variation in the number of hits recorded when different cargoes are stored is too small to be used as a handle to investigate cargo compositions and the degradational impact on the muon tracking for MTS appears to also be very small. These studies are constrained to the impact of secondaries in a simple SilentBorder-like setup, there is significant other work in the realm of using secondary particles to add complementary information to muon tomography techniques. Many of these use specialized separate detector elements to better detect secondary particles [instruments6040066, ; galgoczi2020imaging, ; osti_1097689, ; Bikit_2016, ].
@@ V.1 TomOpt pipeline
TomOpt [TomOpt-Strong2023, ] is a software tool developed for task-specific and constraint-aware optimization of the geometries and design specifications of detectors used for muon scattering tomography. Its modular Python [python, ] architecture features a fully differentiable, end-to-end pipeline that employs differentiable programming through utilizing the automatic differentiation capabilities of the leveraged PyTorch library [pytorch, ]. The optimization pipeline seamlessly integrates the various stages, including muon generation, muon detection and their propagation in scanned volumes leading to volume reconstruction, loss evaluation, and gradient backpropagation with respect to the detector design parameters. The detector parameters are then updated through gradient descent in the direction that minimizes the loss. A schematic representation of this pipeline is shown in FIG. 15, illustrating the main stages in the process, which are further detailed below. The modularity of the simulation chain allows for tailored user-specific implementations of muon source and propagation models, detector parameterization, volume reconstruction methods, loss function definitions, etc.
@@
Prior to optimization, detectors positioned above and below a volume of interest- referred to as the passive volume- are instantiated with sub-optimal parameters. These parameters can be the position in , dimension, efficiency, etc. The passive volume is assigned a voxelized ground truth profile of radiation length (). The optimization then runs over a specified number of epochs, where the detector parameters are updated at the end of each iteration. For each epoch, generated muons are sampled from defined flux models taken from literature [guan2015, , shukla2018, ] and propagated through the detectors and passive volume setup. The scattering model reported by the Particle Data Group [pdg, ] is used for the simulation of the multiple Coulomb scatterings of muons in the passive volume and the calculation of their scattering angles and spatial displacements. The resolution and efficiency of the detectors are modeled by a differentiable surrogate function, represented as a double-sigmoid (FIG. 16). This function achieves maximum resolution and efficiency within the active area of the detector and smoothly decreases while transitioning outside the detector boundaries. Each muon hit in the detectors then is assigned an efficiency and a spatial resolution in which is used to obtain the corresponding reconstructed hit with simulated detector response effect. This is done by sampling uncertainties in and from Gaussian distributions centered at the true position and of standard deviations equal to the corresponding spatial resolutions. Hits outside the detector region are still recorded but with degraded efficiency and resolution, which is a necessary feature to maintain the differentiability of the subsequent volume reconstruction with respect to the detector parameters. This allows for gradient descent to optimize the configuration which records the hits in an optimal way that yields best performance of the studied task. The hits above and below the passive volume are then fitted into incoming and outgoing tracks, respectively, using an analytical likelihood maximization method accounting for the uncertainties on the and positions. The Point-of-Closest-Approach algorithm is a common method used to reconstruct the scattering vertices and obtain a 3D density distribution of these vertices. This approach computes the closest point between the incoming and outgoing muon tracks in 3D space, which corresponds to the midpoint of the common perpendicular to the two tracks. The algorithm assumes that the entirety of the multiple Coulomb scatterings of a muon occurs at this reconstructed point. Subsequent inference methods utilize the spatial and/or angular variables of the reconstructed PoCA points to infer the physical properties of the passive volume under investigation. The discrepancy between the inferred, or predicted, passive volume quantities and their true values is then used to compute a task-specific loss function. Additionally, an optional term reflecting the budget constraints of the detector, calculated based on the unit surface area cost, may be included in the loss function. The loss is then back-propagated through the reconstruction pipeline, and its gradients are calculated with respect to the detector parameters. The detector optimization problem is formulated as a minimization task to be solved through gradient descent. This technique iteratively updates the detector parameters in the direction that minimizes the loss function:
@@ V.2.1 Radiation length estimation
This probabilistic modeling allows the influence of each reconstructed PoCA point to extend across multiple voxels within the passive volume. For each muon batch, each voxel is assigned a probability for every muon’s PoCA point of its being located within that voxel. These probabilities are used as weights to compute a weighted root-mean-square (RMS) of the scattering angle distribution inside each voxel, taking into account all PoCA points in the muon batch. This RMS value is used to estimate the scattering behavior within that voxel. By inverting Eq. 1, the RMS of the scattering angles is employed to infer the radiation length of the material in each voxel. Although the inferred radiation lengths exhibit a systematic bias toward lower values ([TomOpt-Zaher2024, ]), the resulting 3D image effectively distinguishes between high- and low- voxels, providing a clear representation of the material properties within the scanned volume (FIG. 17).
@@ V.2.2 A clustering density inference
Another volume inference method that we adapt in TomOpt is based on a binned clustering algorithm [bca, ] that groups reconstructed scattering vertices, i.e. PoCA points in this case, into clusters indicating different materials in a voxelized volume. This clustering is done based on the spatial proximity and scattering angles of the scattering vertices. Since materials with different densities produce different scattering distributions, the properties of clustered PoCA points can reveal the density profile of the volume and determine high- and low-density regions.
@@
where is the nominal average cosmic muon momentum at sea level. Since most muon tomography setups lack direct momentum measurements for practical reasons, we assume a fixed momentum of 5 GeV for all muons, effectively setting . The median of the natural logarithm of the per voxel weighted metric distribution serves as an indicator of material density, with lower values corresponding to regions where scattering vertices are both larger and closer together. The natural logarithm is applied to improve the separation between closely clustered values. This median value defines the corresponding voxel score, resulting in a per volume distribution of voxel scores. The algorithm subsequently extracts the minimum score value, which serves as a discriminator for identifying the presence of a uranium block in the volume by applying a threshold. While this binary classification approach has been explored in previous work [TomOpt-Zaher2025, ], the present study shifts focus towards the 3D image reconstruction of the passive volume (FIG. 18), utilizing the full voxel score distribution. The previously mentioned extended PoCA pdf in Section V.2.1 is now used, eliminating the need to select a subset of the largest scattering vertices. This allows for a more efficient voxel population, reducing the number of muon events typically required to achieve uniform coverage.
@@ VI.1 The Ax library
While TomOpt’s standard optimization approach relies on gradient descent, we investigate Bayesian Optimization (BO) [Mockus1978, ] as an alternative due to challenges posed by the volume inference methods. In particular, the presence of outliers in the inferred volume properties—especially in the inference approach—leads to a noisy mean squared error (MSE) loss function. Although gradient-based optimization is, in most cases, capable of improving detector coverage by optimizing the spatial positions of the hodoscopes in simple scenarios, such as configurations with a few hodoscopes, the sharp fluctuations in the MSE loss during optimization prevent the final configuration from being considered a fully optimized solution. This issue becomes more pronounced when extending the parameter space to include non-trivial design factors, such as the relative spacing of panels within the hodoscopes.
@@
Bayesian Optimization (BO) is a powerful approach for optimizing black-box functions that are expensive to evaluate, lack a known analytical form or exhibit noise and uncertainty. Unlike gradient-based optimization methods, which require the objective function to be differentiable, BO constructs a probabilistic surrogate model of the objective function that approximates the function and guides the search for the optimal solution. At the core of BO is Bayes’ theorem which states that - given prior knowledge about a function and newly observed data , the posterior distribution of is proportional to the likelihood of given multiplied by the prior distribution of [Brochu2010, ]:
A common choice for the surrogate model in BO is a Gaussian Process (GP), a flexible, non-parametric model that provides a probabilistic estimate of the objective function. A GP assumes that the function values at different points are correlated and defines a prior distribution over possible functions. Given a set of observed data points, the GP updates its belief about the function by computing a posterior distribution, which provides both a mean prediction (expected function value) and a variance estimate (uncertainty in the prediction). This uncertainty measure is particularly advantageous in BO, as it enables informed decisions about where to evaluate the function next. The selection of evaluation points is determined by an acquisition function, which defines a utility function from the model posterior that determines the next point to evaluate [Brochu2010, ]. The selection is performed in a way that strategically balances exploration (sampling uncertain regions) and exploitation (sampling regions with promising performance). One widely used acquisition function is Expected Improvement (EI), which quantifies how much improvement we expect by sampling at a particular point. In the case of minimizing the objective function, the improvement at a new point is defined as the positive difference between the current best observed minimum and the predicted function value , if the new point offers an improvement:
where is the lowest observed function value so far and is the function value at the candidate point . The EI is then computed by taking the expectation of over the posterior distribution of , modeled by the GP [Jones1998, ]:
@@
Ax [Ax, ] is an open-source platform for adaptive experimentation and Bayesian Optimization, developed by Meta. It provides a high-level interface for optimizing parameters without requiring explicit definition of the acquisition function. Ax internally leverages BoTorch [botorch, ], a library for Bayesian Optimization, and automatically selects an acquisition function based on the optimization scenario.
In this work, BO is implemented using the Ax Loop API, which provides an automated framework for sequential BO. The optimization pipeline consists of defining the search space, constructing and updating the surrogate model, and iteratively selecting and evaluating parameter configurations. The search space defines the range of possible values for the parameters being optimized, in our case the positions of hodoscopes. This parameter space is passed to the experiment object, which acts as the central controller for the optimization process. The experiment object in Ax manages the entire optimization process, including the collection of observations, selection of new trials, and updating of the surrogate model. It stores all configurations tested so far along with their corresponding objective function values. Evaluation of each candidate set of parameters is done through a user-defined evaluation function that quantified the performance of the parameters. We take the MSE of the volume inference as our evaluation function. BO with Ax’s Loop API is conducted using a hybrid sampling strategy that consists of two sequential phases:
@@
We use a passive volume with dimensions , simulated as a lorry with an iron base and filled with a random arrangement of iron, beryllium and air blocks. Additionally, a uranium block is randomly placed inside the lorry with a probability of 0.5 of being present. The initial detector setup consists of a hodoscope pair, each measuring , placed above and below the passive volume. Initially, they are offset by in both and directions, leaving part of the passive volume uncovered (FIG. 20).
The reason for placing the hodoscopes at different positions is to avoid overlap along the axis. This is due to the current implementation of hodoscopes in TomOpt, where the detector plates within each hodoscope are defined as unique plates. The fitting of hits into tracks accounts for hits in all detector plates, and placing two plates at the same level would cause a problem. This limitation has been addressed in a new implementation, which is currently being studied for a future paper. Starting from this suboptimal configuration, the optimization is expected to converge toward an optimal final setup where the hodoscopes perfectly cover the passive volume. BO is then executed for evaluation trials. Before each evaluation, the detector parameters are set to the values sampled for evaluation.

### v2

@@ IV GEANT4 Study of Secondary Hits
As muons pass through the hodoscope, cargo, container, and surrounding air they can produce secondary particles, primarily delta rays. An assessment is made in this section of the impact of any hodoscope hits induced by such secondary activity. Note that activity due to other particles generated in cosmic air showers is not considered in this study but can also be considered as a source of ‘secondary’ information in MST.
@@
The goal of this study was to evaluate what impact deposits from secondaries could have in degrading the performance of MST with the SilentBorder scanners. It is clear from Fig. 6 that above a certain energy threshold there is no method to distinguish between secondary and muon hits based on the scale of the energy deposition.
For the first study a threshold of 0.4 MeV is applied, retaining only those depositions that lie within the MIP peak region. A cargo block of size 1 m1 m0.5 m is placed at the center of the container with a variety of material types considered. The MuSiBo generator is used to provide muons corresponding to a 15 minute exposure time. Fig. 7(a) shows the total number of hits recorded across all hodoscopes per event, regardless of whether they were produced via secondaries or muons. In general the peak structure around 6 hits is produced via the 3 hits left by the muon, before and after the cargo whilst the small tail is produced by the presence of any secondary hits. Whilst it appears that there is some variation between the different cargo types this variation stems from the number of events passing the filter, not from a significant difference in the number of secondary hits produced per muon. Higher density and higher-Z materials reduce the number of events that pass the filtering process. Fig. 7(b) shows the same distribution but normalized by the number of events that passed the filtering, very limited variation per cargo is now observed.
@@
In conclusion, it is clear that with the current SilentBorder scanner setup the impact of secondary hits is very limited. The variation in the number of hits recorded when different cargoes are stored is too small to be used as a handle to investigate cargo compositions and the degradational impact on the muon tracking for MTS appears to also be very small. These studies are constrained to the impact of secondaries in a simple SilentBorder-like setup, there is significant other work in the realm of using secondary particles to add complementary information to muon tomography techniques. Many of these use specialized separate detector elements to better detect secondary particles [Pérez Prada et al., 2022; Galgóczi et al., 2020; Bacon et al., 2013; Bikit et al., 2016].
@@ V.1 TomOpt pipeline
TomOpt [Strong and others, 2024] is a software tool developed for task-specific and constraint-aware optimization of the geometries and design specifications of detectors used for muon scattering tomography. Its modular Python [Python Software Foundation, ] architecture features a fully differentiable, end-to-end pipeline that employs differentiable programming through utilizing the automatic differentiation capabilities of the leveraged PyTorch library [Paszke and others, 2019]. The optimization pipeline seamlessly integrates the various stages, including muon generation, muon detection and their propagation in scanned volumes leading to volume reconstruction, loss evaluation, and gradient backpropagation with respect to the detector design parameters. The detector parameters are then updated through gradient descent in the direction that minimizes the loss. A schematic representation of this pipeline is shown in Fig. 14, illustrating the main stages in the process, which are further detailed below. The modularity of the simulation chain allows for tailored user-specific implementations of muon source and propagation models, detector parameterization, volume reconstruction methods, loss function definitions, etc.
@@
Prior to optimization, detectors positioned above and below a volume of interest- referred to as the passive volume- are instantiated with sub-optimal parameters. These parameters can be the position in , dimension, efficiency, etc. The passive volume is assigned a voxelized ground truth profile of radiation length (). The optimization then runs over a specified number of epochs, where the detector parameters are updated at the end of each iteration. For each epoch, generated muons are sampled from defined flux models taken from literature [Guan et al., 2015, Shukla and Sankrith, 2018] and propagated through the detectors and passive volume setup. The scattering model reported by the Particle Data Group [Navas and others, 2024] is used for the simulation of the multiple Coulomb scatterings of muons in the passive volume and the calculation of their scattering angles and spatial displacements. The resolution and efficiency of the detectors are modeled by a differentiable surrogate function, represented as a double-sigmoid (Fig. 15). This function achieves maximum resolution and efficiency within the active area of the detector and smoothly decreases while transitioning outside the detector boundaries. Each muon hit in the detectors then is assigned an efficiency and a spatial resolution in which is used to obtain the corresponding reconstructed hit with simulated detector response effect. This is done by sampling uncertainties in and from Gaussian distributions centered at the true position and of standard deviations equal to the corresponding spatial resolutions. Hits outside the detector region are still recorded but with degraded efficiency and resolution, which is a necessary feature to maintain the differentiability of the subsequent volume reconstruction with respect to the detector parameters. This allows for gradient descent to optimize the configuration which records the hits in an optimal way that yields best performance of the studied task. The hits above and below the passive volume are then fitted into incoming and outgoing tracks, respectively, using an analytical likelihood maximization method accounting for the uncertainties on the and positions. The Point-of-Closest-Approach algorithm is a common method used to reconstruct the scattering vertices and obtain a 3D density distribution of these vertices. This approach computes the closest point between the incoming and outgoing muon tracks in 3D space, which corresponds to the midpoint of the common perpendicular to the two tracks. The algorithm assumes that the entirety of the multiple Coulomb scatterings of a muon occurs at this reconstructed point. Subsequent inference methods utilize the spatial and/or angular variables of the reconstructed PoCA points to infer the physical properties of the passive volume under investigation. The discrepancy between the inferred, or predicted, passive volume quantities and their true values is then used to compute a task-specific loss function. Additionally, an optional term reflecting the budget constraints of the detector, calculated based on the unit surface area cost, may be included in the loss function. The loss is then back-propagated through the reconstruction pipeline, and its gradients are calculated with respect to the detector parameters. The detector optimization problem is formulated as a minimization task to be solved through gradient descent. This technique iteratively updates the detector parameters in the direction that minimizes the loss function:
@@ V.2.1 Radiation length estimation
The probabilistic modeling of PoCA vertices allows the influence of each reconstructed PoCA point to extend across multiple voxels within the passive volume. For each muon batch, each voxel is assigned a probability for every muon’s PoCA point of its being located within that voxel. These probabilities are used as weights to compute a weighted root-mean-square (RMS) of the scattering angle distribution inside each voxel, taking into account all PoCA points in the muon batch. This RMS value is used to estimate the scattering behavior within that voxel. By inverting Eq. 1, the RMS of the scattering angles is employed to infer the radiation length of the material in each voxel. Although the inferred radiation lengths exhibit a systematic bias toward lower values ([Zaher and others, 2024]), the resulting 3D image effectively distinguishes between high- and low- voxels, providing a clear representation of the material properties within the scanned volume (Fig. 16).
@@ V.2.2 A clustering density inference
Another volume inference method that we adopt in TomOpt is based on a binned clustering algorithm [Thomay and others, 2013] that groups reconstructed scattering vertices, i.e. PoCA points in this case, into clusters indicating different materials in a voxelized volume. This clustering is done based on the spatial proximity and scattering angles of the scattering vertices. Since materials with different densities produce different scattering distributions, the properties of clustered PoCA points can reveal the density profile of the volume and determine high- and low-density regions.
@@
where is the nominal average cosmic muon momentum at sea level. Since most muon tomography setups lack direct momentum measurements for practical reasons, we assume a fixed momentum of 5 GeV for all muons, effectively setting . The median of the natural logarithm of the per voxel weighted metric distribution serves as an indicator of material density, with lower values corresponding to regions where scattering vertices are both larger and closer together. The natural logarithm is applied to improve the separation between closely clustered values. This median value defines the corresponding voxel score, resulting in a per volume distribution of voxel scores. The algorithm subsequently extracts the minimum score value, which serves as a discriminator for identifying the presence of a uranium block in the volume by applying a threshold. While this binary classification approach has been explored in previous work [Zaher et al., 2025], the present study shifts focus towards the 3D image reconstruction of the passive volume (Fig. 17), utilizing the full voxel score distribution. The previously mentioned extended PoCA pdf in Section V.2.1 is now used, eliminating the need to select a subset of the largest scattering vertices. This allows for a more efficient voxel population, reducing the number of muon events typically required to achieve uniform coverage.
@@ VI.1 The Ax library
While TomOpt’s standard optimization approach relies on gradient descent, we investigate Bayesian Optimization (BO) [Mockus et al., 1978] as an alternative due to challenges posed by the volume inference methods. In particular, the presence of outliers in the inferred volume properties—especially in the inference approach—leads to a noisy mean squared error (MSE) loss function. Although gradient-based optimization is, in most cases, capable of improving detector coverage by optimizing the spatial positions of the hodoscopes in simple scenarios, such as configurations with a few hodoscopes, the sharp fluctuations in the MSE loss during optimization prevent the final configuration from being considered a fully optimized solution. This issue becomes more pronounced when extending the parameter space to include non-trivial design factors, such as the relative spacing of panels within the hodoscopes.
@@
Bayesian Optimization (BO) is a powerful approach for optimizing black-box functions that are expensive to evaluate, lack a known analytical form or exhibit noise and uncertainty. Unlike gradient-based optimization methods, which require the objective function to be differentiable, BO constructs a probabilistic surrogate model of the objective function that approximates the function and guides the search for the optimal solution. At the core of BO is Bayes’ theorem which states that - given prior knowledge about a function and newly observed data , the posterior distribution of is proportional to the likelihood of given multiplied by the prior distribution of [Brochu et al., 2010]:
A common choice for the surrogate model in BO is a Gaussian Process (GP), a flexible, non-parametric model that provides a probabilistic estimate of the objective function. A GP assumes that the function values at different points are correlated and defines a prior distribution over possible functions. Given a set of observed data points, the GP updates its belief about the function by computing a posterior distribution, which provides both a mean prediction (expected function value) and a variance estimate (uncertainty in the prediction). This uncertainty measure is particularly advantageous in BO, as it enables informed decisions about where to evaluate the function next. The selection of evaluation points is determined by an acquisition function, which defines a utility function from the model posterior that determines the next point to evaluate [Brochu et al., 2010]. The selection is performed in a way that strategically balances exploration (sampling uncertain regions) and exploitation (sampling regions with promising performance). One widely used acquisition function is Expected Improvement (EI), which quantifies how much improvement we expect by sampling at a particular point. In the case of minimizing the objective function, the improvement at a new point is defined as the positive difference between the current best observed minimum and the predicted function value , if the new point offers an improvement:
where is the lowest observed function value so far and is the function value at the candidate point . The EI is then computed by taking the expectation of over the posterior distribution of , modeled by the GP [Jones et al., 1998]:
@@
Ax [Meta Open Source, ] is an open-source platform for adaptive experimentation and Bayesian Optimization, developed by Meta. It provides a high-level interface for optimizing parameters without requiring explicit definition of the acquisition function. Ax internally leverages BoTorch [Balandat et al., 2020], a library for Bayesian Optimization, and automatically selects an acquisition function based on the optimization scenario.
In this work, BO is implemented using the Ax Loop API, which provides an automated framework for sequential BO. The optimization pipeline consists of defining the search space, constructing and updating the surrogate model, and iteratively selecting and evaluating parameter configurations. The search space defines the range of possible values for the parameters being optimized, in our case the positions of hodoscopes. This parameter space is passed to the experiment object, which acts as the central controller for the optimization process. The experiment object in Ax manages the entire optimization process, including the collection of observations, selection of new trials, and updating of the surrogate model. It stores all configurations tested so far along with their corresponding objective function values. Evaluation of each candidate set of parameters is done through a user-defined evaluation function that quantified the performance of the parameters. We take the MSE of the volume inference as our evaluation function, which quantifies the accuracy of the voxelized reconstructed material distribution and directly relates to the ability to discriminate high-density or suspicious objects in border security applications. BO with Ax’s Loop API is conducted using a hybrid sampling strategy that consists of two sequential phases:
@@
[…]
Initially, hodoscopes are offset by in both and directions, leaving part of the passive volume uncovered (Fig. 19).
The reason for placing the hodoscopes at different positions is to avoid overlap along the axis. This is due to the current implementation of hodoscopes in TomOpt, where each hodoscope is defined as a single unit consisting of individual detector plates. The fitting of hits into tracks accounts for hits in all detector plates, where hits outside an active plate surface are down-weighted through lower resolution and efficiency. Hence, placing two plates at the same level would cause ambiguities in track reconstruction. While this constraint does not prevent the optimisation study focusig on inference methods, it does limit the ability to model perfectly co-planar panels or more complex detector layouts where modules intentionally share the same z-position.This panel overlap limitation is a technical constraint of the current version and is being addressed in an updated implementation that will be presented in a forthcoming paper.
Starting from this suboptimal configuration, the optimization is expected to converge toward an optimal final setup where the hodoscopes perfectly cover the passive volume. BO is then executed for evaluation trials. Before each evaluation, the detector parameters are set to the values sampled for evaluation.